A coffee counter charges $9.00 per pound for french roast, $8.00 a pound for sumantra. How much of each type should be used to make a 20 pound blend that sells for $8.65 per pound?
Answer by mananth(16946)

---------price ---------------- quantity
French roast 9 ---------------- x pounds
sumantra 8 ------ 20 - x pounds
Mixture 8.65% ---------------- 20

9 x + 8 ( 20 - x ) = 8.65 * 20

9x+160-8x=173
9x-8x=173-160
x = 13
x = 13pounds $9.00 French roast
7 pounds $8.00 sumantra